Yesteryear Productions is considering a project with an initial start up cost of $740,000. The firm maintains a debt-equity ratio of .50 and has a flotation cost of debt of 6.2 percent and a flotation cost of equity of 13.1 percent. The firm has sufficient internally generated equity to cover the equity cost of this project. What is the initial cost of the project including the flotation costs?
Answer #1

Let Equity be 1 ,so Debt = 1*.50 = .50 , so weight of Debt = .50/(1+.50) = .33333 and weight of equity = 1-.33333= .66667

Weighted average Flotation cost =[Flotation cost for debt *weight of debt]+[Flotation cost for equity * weight of equity]

         =[6.2 * .33333]+[0 * .66667]

        = 2.06665+ 0

       = 2.0667%

***No flotation cost for equity since firm has sufficient funds to cover for equity .

Initial cost = Initial start up cost /(1-F)

        = 740000/(1-.020667)

        =740000 / .97933

       = $ 755618